# The unified cross-disciplinary model of the operation of neurons

### v2

@@ 4.2 The oscillator model
Despite the early warning that ’it was not possible to separate the change into resistance and capacity components[31], a commonly accepted truism was that neurons, in some sense, behave as electric oscillators. HH introduced the idea explicitly that the electrically equivalent circuit of a neuron is an $RC$ oscillator. We must also add the invention from about three decades later: "Neurons ensure the directional propagation of signals throughout the nervous system. The functional asymmetry of neurons is supported by cellular compartmentation: the cell body and dendrites (somatodendritic compartment) receive synaptic inputs, and the axon propagates the action potentials that trigger synaptic release toward target cells. Between the cell body and the axon sits a unique compartment called the axon initial segment. The AIS was first described 50 years ago [i.e., nearly two decades after HH published their study], and its molecular composition and organization have been progressively elucidated during the following decades. …. Recent years have also brought crucial insights into the functions of the AIS: how ion channels at its surface generate and shape the action potential." [12] We provide the physics and mathematics of how AIS shapes the action potential (or more precisely, we show what an important role it plays in forming AP).
Inventing AIS changed the viewpoint of neuroscience [12], a half century later, after setting up the commonly accepted model by HH. "The axon initial segment (AIS) is located at the proximal axon and is the site of action potential initiation. This reflects the high density of ion channels found at the AIS. … The summation of synaptic inputs gives rise to action potentials at the axon initial segment (AIS), a 20–60 $\mu m$ long domain located at the proximal axon/soma interface that has a high density of voltage-gated ion channels." As discussed in [13], see also their Figure 1, the structure of the Axon Initial Segment is known to the most minor details. As the illuminating investigations in 2008 [10] revealed, the AIS has very dense ion channels. That is, from an electrical point of view, those parallelized channels can be abstracted as a discrete conductance (or resistance) between the membrane and the axon. The membrane itself can be abstracted as a distributed condenser with no resistance. This model is in contrast with the viewpoint of biophysics, that the membrane plus AIS is considered a distributed element, where the capacitor and condenser cannot be separated. The current model is the result of projecting the physical processes in the resting state to the transient state. Notice the important point: "Neurons are also anatomically polarized, as they can be subdivided into a somatodendritic input domain and an axonal output domain" [9]; providing direct evidence that (unlike in HH’s model) the input and output currents (and voltage time derivatives) are independent.
@@
In electric view, the process can be modeled as follows: a sudden voltage gradient (see Figure 12 middle inset) appears on the membrane (as a discrete capacitor), and a current flows out (see Figure 12 bottom inset) through the serially connected ion channel array (as a discrete resistor). The neuron forms a simple serially (not parallelly, as assumed by Hodgkin and Huxley [49] and mistakenly claimed by neurophysiology) connected $RC$ oscillator. For the physical and mathematical description, see section 4.3, furthermore, [63], section 4; for its algorithmic specifics [64], for further details [16].
Figure 12: The physical processes describing the membrane’s operation. The rushed-in $Na^{+}$ ions instantly increase the charge on the membrane, then the membrane’s capacity discharges (produces an exponentially decaying voltage derivative). The ions are created at different positions on the membrane, so they need different times to reach the AIS, where the current produces a peaking voltage derivative. The resulting voltage derivative, the sum of the two derivatives (the AIS current is outward), drives the oscillator. Its integration produces the membrane potential. When the membrane potential crosses the voltage threshold value, it switches the synaptic currents off/on.
The fundamental difference between the parallel and serial $RC$ circuits is that the parallel circuit cannot produce output voltage with opposite sign. As we discussed, the capacitive current, which by definitionem changes its direction and so generates an opposite voltage on the resistor represented by the AIS, perfectly describes the so-called "hyperpolarization". It is not the effect of a $K^{+}$ current: the resting ion channels do not have enough transmission capacity. Not knowing about the AIS leads to assuming that the resting ion channels work also as transient channels (in other words, assuming a parallel $RC$ circuit instead of the serial one, furthermore, that the output current flows through the membrane instead of the axon). Some $K^{+}$ current certainly exists; for the magnitude of currents during AP, see [15].
The shape of the output waveform depends on the pulse width ratio to the $RC$ time constant. When $RC$ is much larger than the pulse width, the output waveform resembles the input signal, even with a square wave input. (In the case of the neuronal oscillator, the shape of the front side of the spike is similar to the one derived for the rush-in current, while the back side is very prolonged.)
@@ 4.3 Mathematics of the ActionPotential
Although the implementation of a neuronal circuit works with ionic currents and has a charge transmission mechanism drastically different from the one based on free electrons used in metals, the operation can align (to some measure) with the instant currents of physical electronics (as we suggest, the effect of the slow current and finite size can be reasonably imitated with current generators generating a special current shape, see section 2.10). The input currents in a biological circuit are a large rush-in current through the membrane in an extremely short time, and much smaller gated synaptic currents arriving at different times through the synapses. Those currents generate voltage gradients, and those gradients direct the operation of the $RC$ oscillator (Notice that due to the multi-physical effects, the parameters $R$ and $C$ may change in the function of time and the actual voltage). As it is well known from the theory of electric circuits, the output voltage measured on the output resistor is as follows:
@@
that is, the (temporally gated) sum of the input gradient that the currents generate, plus the gradient of the output current through the AIS [13, 12]. The latter term can be described as follows:
$\frac{d}{dt}V_{OUT}^{AIS}={\frac{1}{C}}\frac{V_{internal}-V_{external}}{R}$ (42)
The input currents are slow (we typically imitate them by a current generator with a special form), see section 2.10. In the simplest case, the resulting voltage derivative comprises only the rush-in contribution described by Equation (32) and the AIS contribution described by Equation (42). The gating implements a dynamically changing temporal window and dynamically changing synaptic weights. The appearance of the first arriving synaptic gradient starts the “Computing” stage. The rush-in gradient starts the “Delivering” stage [64]. The charge collected in that temporal window is as follows:
@@ 4.4 Damped oscillation
A considerable resemblance with the shape of an AP is evident, but there exist considerable differences from the ’net’ physical case. Among others, the mass $m$ changes continuously; furthermore, it is distributed in the volume of the neuron, so one cannot expect a perfect description from such a simple motion equation. However, it is perfect for demonstrating that such a system can produce the symptoms of relaxation and hyperpolarization that physiology observes. (The electrostatic interaction is the strongest and fastest interaction in nature, stronger than the other interactions that have a role in the process, so we have reasons to think that Eq. (40) provides the best description.)
The "moving force" has two major components. The sudden rush-in of $Na^{+}$ ions has (at least) a double effect. One effect is that they press the surface of the membrane that elastically increases its diameter [70]. Due to their mutual repulsion, the ions experience a strong electric force toward the membrane. It represents an impulse $J=F\,\Delta t$ (N  s), that enables estimating the time and energy of the change the rush-in causes and explains that the energy needed to generate an AP is suddenly produced electrically and stored temporarily as elastic energy (the energy is produced in the background, when the ATP produces ions by hydrolysis, as we described above); see Appendix A. At the same time, it offers a plausible way of unifying the electric and thermodynamical attempts to describe neuronal operation; furthermore, a pictorial understanding of why the elastic change produces opposite amplitudes in the size of the membrane and the voltage in its surface layer. Similarly, it explains that in the first phase of generating an AP the neuron invests (performs a work $Q_{H}$ against the membrane’s elastic force), while in the second phase the electic membrane performs work $Q_{L}$ on pressing out the ions physiology observed in an AP; for the details see section 4.5. These two changes are inseparable, and their combined effect can explain the observed consequences. The compression and expansion give way to explaining the observed mechanical, optical, thermal, etc. changes.
@@
Their other effect is that the appearance of the ions drastically increases the potential in the surface layer of the neuron membrane. During regular operation, a very intense current pulse is created: "A small flow of ions carries sufficient charge to cause a large change in the membrane potential. The ions that give rise to the membrane potential lie in a thin (<1nm<1\ nm) surface layer close to the membrane" [45] (see the dotted line in Fig. 4; notice that the horizontal scale on the figure spans a few $nm$). For this, a considerable number of ion channels are used. When the trigger signal is received, the gated ion channels in the membrane’s wall open quickly and only for a brief period. The rushed-in ions create a large offset potential in the layer proximate to the membrane’s surface (from an electrical point of view, a condenser) that drives an intense current through the always-open ion channel array AIS (from an electrical point of view, a resistor) at the beginning of the axon. The slow current on the membrane’s surface mimics a condenser: in the picture of classical electricity, the charges stay inside the distributed condenser while they travel on its surface. In classical electronics, the condenser is sizeless.
4.5 Thermodynamic aspects
As the middle inset in Fig. 12 shows, the sum of the rush-in and AIS gradients controls generating the AP. During excitation, $Q_{X}$ energy is fed into the neuron. The excitation increases the neuron’s energy by $Q_{H}$. When the membrane potential reaches its peak value, the gradient changes its sign and starts to decrease the potential. In the transient state, the size of the neuron (due to the increased electrostatic repulsion) slightly increases [70] and stores the energy mainly in the form of elastic energy. The resultant of the electric, elastic and thermodynamic forces moves the ions toward the AIS. Although part of the energy is dissipated while the ions exit through the AIS with their Stokes-Einstein speed, the rest of the energy (the elastic part) is recovered. The concentration and potential values change as discussed in connection with Fig. 10. The membrane voltage controls synapses as shown in the top inset in Fig. 12.
@@ 4.5 Thermodynamic aspects
Figure 15: The time course of the electric gradient [54] perfectly matches that of the mechanical and thermal changes [27]; however, classic theory cannot explain the connection between them [8]. The experimental $\frac{dV}{dt}$ uses an opposite sign convention.
@@
The ‘phase loop’ (how the voltage gradient and membrane voltage correlate; a parametric curve that uses time $t$ as parameter) shown on the left side of Fig. 14 represents a strong test of the theoretical line shape: the values of the potential and its gradient are simultaneously compared to the experimentally measured ones. Comparing the experimental and theoretical shapes supports the theory that describes the mutual dependence of voltage and its gradient on each other. Furthermore, it also shows that omitting the gradient from the classic theory misled the experiment designers, and the measurement design was wrong: the time interval between recording data points was too long. Where the gradient is high and changes quickly (see Fig. 12, middle inset), the blue experimental diagram line is roughly broken. The changes in the values are too significant, causing the integration to distort the value and resulting in an inaccurate broken line. In the rush-in region, the theoretical line is even out of range of the figure. Where the gradient change is slow, the measurement points are sufficiently dense, and the theory perfectly describes the experimental data. The classical theories cannot interpret the phenomenon and do not produce any similar dependence; furthermore, they affect the experiment design in the wrong direction.
@@
An exciting aspect is that the phase plot in Fig. 14, left side, shows a remarkable resemblance to a specialized thermodynamic Carnot cycle in Fig. 14, right side; and inspires us to interpret the parameters of the "neuronal Carnot machine" (the bent arcs serve only for illustration; they do not follow the path on the left side). As discussed in [27], a neuron emits and absorbs heat (energy) in the first and second phases, respectively, of an AP; see Fig. 15. The experience, alone, defies the classical HH mechanism. Laws describing currents of electrons must not be used directly to currents of ions. Neglecting the repelling force between the charge carriers of the ion current, which can be omitted in the case of electron current, leads to entirely wrong conclusions, from the wrong oscillator model to losing the connection between the electrical and mechanical changes, to losing the connection between inanimate and living nature. Our theoretical model, unlike all the other models, can explain that observation quantitatively.
@@
We rescale the horizontal axis with the resistance $R$ of the AIS (actually, the horizontal axis is transformed to $I_{AIS}$); thus, the plot becomes a neuronal energy density diagram. Actually, it corresponds to a usual PVP-V diagram, and we expect that the temperature changes in the same analogy (we keep the convention of using the letter $V$ to denote both potential and volume). By integrating the parametric $t$ values of any two points on the diagram ‘circle’ line, we receive the work performed by the neuron.
@@
We consider that the environment in the period ‘Excite’ invests energy $Q_{X}$ through the neuron’s synapses (increases the membrane’s voltage to the threshold) between [0,1] to open the valves. Opening the valves enables the neuron receiving in period ‘Rushin’ energy $Q_{R}$ by letting in $Na^{+}$ ions, accelerated by the field across the membrane. Suddenly, the ion concentration and the membrane potential increase to their maximum value. Given that the process is instant, the potential $V$ does not change (no current flows), so the integral [1,2] evaluates to zero (an ‘iso-voltage’ transition). Now the neuron has a high potential to prepare its (already encoded) message (the AP). As our model explains, in the first phase (from the threshold potential to the maximum value of the AP, where $\frac{dV}{dt}$ is positive), the neuronal power cycle produces a positive heat $Q_{H}$, given that $\frac{dV}{dt}$ is always positive. In the second phase (from the top value of AP to where the driving force $\frac{dV}{dt}$ becomes zero again), the neuronal power cycle produces a negative heat $Q_{L}$, given that $\frac{dV}{dt}$ is always negative; in line with the measurement results shown in Fig. 15 [27]. The integral [2,3] provides the “heating” $Q_{H}$, the energy invested. Similarly, the integral [3,6] provides the “cooling” $Q_{L}$, representing the energy recovered (assuming no artificial current). In the “Relax” phase, $\frac{dV}{dt}$ is (almost) zero, so the corresponding integral evaluates to zero (‘iso-gradient’ transition). In this phase, the neuron must restore its resting state: by using ATP, it creates new ions. The difference in the energies of $Q_{H}$ and $Q_{L}$ is the net energy used to forward an AP; while $Q_{X}$ is the energy used to perform the computation. The energy $Q_{X}$ is an approximate value: as shown in Fig. 12, at the time of re-opening the synaptic inputs, the membrane can be above or below the resting potential, which actually means a neuron-level memory. In this sense, adjacent neural spikes can borrow energy from each other (this way explaining why the first spikes in a burst behave differently). The energy $Q_{R}$ is produced by the ATP hydrolysis in parallel with the ’primary processes’ described here, and the produced ions are acquired on the condenser plates, thus restoring the resting potential and providing potential energy for the forthcoming spikes.
@@
From Fig. 14, after recalibration (with $R_{AIS}=10^{7}$ Ohm), one can estimate (by approximating roughly the area under the arcs by triangles, in units of $10^{-7}[J]$), $Q_{H}=5.2,Q_{L}=1.8,W=3.4,Q_{X}=0.2$. That is, according to the "phase plot" [54], the neuron operates with an efficiency around 65%, and approximately 3% of the energy is derived from the upstream neurons. The values (without using data from dedicated measurements) agree excellently with the measured [15] values $W=2.5$ and 74%; furthermore, the $Q_{X}$:total ratio $1:35$, derived by measuring the direct energy consumption [14].
In our model, the $Na^{+}$ rush-in increases the potential, the concentration, the pressure (due to electric repulsion; so it changes the membrane’s capacity), and the volume (since the membrane is elastic), so a multi-dimensional Carnot diagram could describe the state accurately (although the invested energy can be safely estimated from the work the charge performs given that the energy is almost entirely stored as elastic potential energy due to the enormous elasticity modulus). Furthermore, all parameters have a time course. It is not possible to provide a more accurate estimation without a true multi-physics simulation. However, the approach is promising.

### v3

@@ 6.2 The oscillator model
Despite the early warning that ’it was not possible to separate the change into resistance and capacity components[50], a commonly accepted truism was that neurons, in some sense, behave as electrical oscillators. HH introduced the idea explicitly that the electrically equivalent circuit of a neuron is an $RC$ oscillator. We must also add the invention from about three decades later: "Neurons ensure the directional propagation of signals throughout the nervous system. The functional asymmetry of neurons is supported by cellular compartmentation: the cell body and dendrites (somatodendritic compartment) receive synaptic inputs, and the axon propagates the action potentials that trigger synaptic release toward target cells. Between the cell body and the axon sits a unique compartment called the axon initial segment. The AIS was first described 50 years ago [i.e., nearly two decades after HH published their study], and its molecular composition and organization have been progressively elucidated during the following decades. …. Recent years have also brought crucial insights into the functions of the AIS: how ion channels at its surface generate and shape the action potential." [21] We provide the physics and mathematics of how AIS shapes the action potential (or more precisely, we show what an important role it plays in forming AP).
Inventing AIS changed the viewpoint of neuroscience [21], a half century later, after setting up the commonly accepted model by HH. "The AIS is located at the proximal axon and is the site of action potential initiation. This reflects the high density of ion channels found at the AIS. … The summation of synaptic inputs gives rise to action potentials at the AIS, a 20–60 $\mu m$ long domain located at the proximal axon/soma interface that has a high density of voltage-gated ion channels." As discussed in [71], see also their Figure 1, the structure of the AIS is known to the most minor details. As the illuminating investigations in 2008 [20] revealed, the AIS has very dense ion channels. That is, from an electrical point of view, those parallelized channels can be abstracted as a discrete conductance (or resistance) between the membrane and the axon. The membrane itself can be abstracted as a distributed condenser with no resistance. This model is in contrast with the viewpoint of biophysics, that the membrane plus AIS is considered a distributed element, where the capacitor and condenser cannot be separated. The current model is the result of projecting the physical processes in the resting state to the transient state. Notice the important point: "Neurons are also anatomically polarized, as they can be subdivided into a somatodendritic input domain and an axonal output domain" [69]; providing direct evidence that (unlike in HH’s model) the input and output currents (and voltage time derivatives) are independent.
@@
In electrical view, the process can be modeled as follows: a sudden voltage gradient (see Figure 13 middle inset) appears on the membrane (as a discrete capacitor), and a current flows out (see Figure 13 bottom inset) through the serially connected ion channel array (as a discrete resistor). The neuron forms a simple serially (not parallelly, as assumed by Hodgkin and Huxley [14] and mistakenly claimed by neurophysiology) connected $RC$ oscillator. For the physical and mathematical description, see section 6.3, furthermore,  section 4 in [73]; for its algorithmic specifics [5], for further details [72]. Notice that the difference between the rush-in and the resultant gradients emphasizes the role of the finite resources: without the AIS, no turn-back of the resultant gradient would occur.
Figure 13: The physical processes describing the membrane’s operation. The rush-in $Na^{+}$ ions instantly increase the membrane’s charge, and the membrane’s capacity discharges (producing an exponentially decaying voltage derivative). The ions are created at different positions on the membrane, so they take different times to reach the AIS, where the current produces a peak in the voltage derivative. The resulting voltage derivative, the sum of the two derivatives (the AIS current is outward), drives the oscillator. Its integration produces the membrane potential. When the membrane potential crosses the threshold, it switches the synaptic currents on or off.
The fundamental difference between the parallel and serial $RC$ circuits is that the parallel circuit cannot produce output voltage with opposite sign. As discussed, the capacitive current, which by definition changes its direction and so generates an opposite voltage on the resistor represented by the AIS, perfectly describes the so-called "hyperpolarization". It is not the effect of a $K^{+}$ current: the resting ion channels do not have sufficient conductance. Not knowing about the AIS leads to assuming that the resting ion channels work also as transient channels (in other words, assuming a parallel $RC$ circuit instead of the serial one, furthermore, that the output current flows through the membrane instead of the axon). Some $K^{+}$ current certainly exists; for the magnitude of currents during AP, see [43].
The shape of the output waveform depends on the ratio of the pulse width to the $RC$ time constant. When $RC$ is much larger than the pulse width, the output waveform resembles the input signal, even with a square wave input. (In the case of the neuronal oscillator, the shape of the front side of the spike is similar to the one derived for the rush-in current, while the back side is very prolonged.)
@@ 6.3 Mathematics of the ActionPotential
Although the implementation of a neuronal circuit works with ionic currents and has a charge transmission mechanism drastically different from the one based on free electrons used in metals, the operation can align (to some measure) with the instant currents of physical electronics (as we suggest, the effect of the slow current and finite size can be reasonably imitated with current generators generating a special current shape, see section 4.9). The input currents in a biological circuit are a large rush-in current through the membrane in an extremely short time, and much smaller gated synaptic currents arriving at different times through the synapses. Those currents generate voltage gradients, and those gradients direct the operation of the $RC$ oscillator (Notice that due to the multi-physical effects, the parameters $R$ and $C$ may change in the function of time and the actual voltage). As it is well known from the theory of electrical circuits, the output voltage measured on the output resistor is as follows:
@@
that is, the (temporally gated) sum of the input gradient that the currents generate, plus the gradient of the output current through the AIS [71, 21]. The latter term can be described as follows (see the discussion around also Eq.(43):
$\frac{d}{dt}V_{OUT}^{AIS}={\frac{1}{C}}\frac{V_{internal}-V_{external}}{R_{AIS}}$ (48)
The input currents are slow (we typically imitate them using a current generator with a special form); see section 4.9. In the simplest case, the resulting voltage derivative comprises only the rush-in contribution described by Equation (35) and the AIS contribution described by Equation (48). The gating implements a dynamically changing temporal window and dynamically changing synaptic weights. The appearance of the first arriving synaptic gradient starts the “Computing” stage. The rush-in gradient starts the “Delivering” stage [5]. The charge collected in that temporal window is as follows:
@@ 6.4 Damped oscillation
A remarkable resemblance with the shape of an AP is evident, but there exist significant differences from the ’net’ physical case. Among others, the mass $m$ changes continuously; furthermore, it is distributed in the volume of the neuron, so one cannot expect a perfect description from such a simple motion equation. However, it is perfect for demonstrating that such a system can produce the symptoms of relaxation and hyperpolarization that physiology observes. (The electrostatic interaction is the strongest and fastest interaction in nature, stronger than the other interactions that have a role in the process, so we have reasons to think that Eq. (46) provides the best description.)
The "moving force" has two major components. The sudden rush-in of $Na^{+}$ ions has (at least) a double effect. One effect is that they press the surface of the membrane that elastically increases its diameter [80]. Due to their mutual repulsion, the ions experience a strong electrical force toward the membrane. It represents an impulse $J=F\,\Delta t$ (N  s), that enables estimating the time and energy of the change the rush-in causes and explains that the energy needed to generate an AP is suddenly produced electrically, and stored temporarily as elastic energy (the neuron produces the required energy in the background, when ATP produces ions by hydrolysis, as described above); see Appendix B. The case is practically identical to the deformation of an elastic plate induced by the hydrostatic pressure of a water column. At the beginning, the aluminum plate is suddenly subjected to hydrostatic pressure, and it reaches equilibrium after initial oscillations. The diagram line of the process is shown in the figure [81], and the method of simulation is discussed in [82]. Compare the diagram line to the gradient in Fig. 13, middle inset.
At the same time, it offers a plausible way of unifying the electrical and thermodynamic attempts to describe neuronal operation; furthermore, a pictorial understanding of why the elastic change produces opposite amplitudes in the size of the membrane and the voltage in its surface layer. Similarly, it explains that in the first phase of generating an AP the neuron invests (performs a work $Q_{H}$ against the membrane’s elastic force), while in the second phase the electic membrane performs work $Q_{L}$ on pressing out the ions physiology observed in an AP; for the details see section 7. These two changes are inseparable, and their combined effect can explain the observed consequences. The compression and expansion give way to explaining the observed mechanical, optical, thermal, etc. changes.
@@
Their other effect is that the appearance of the ions drastically increases the potential in the surface layer of the neuron membrane. During regular operation, a very intense current pulse is created: "A small flow of ions carries sufficient charge to cause a large change in the membrane potential. The ions that give rise to the membrane potential lie in a thin (<1nm<1\ nm) surface layer close to the membrane" [41] (see the dotted line in Fig. 4; notice that the horizontal scale on the figure spans a few $nm$). For this, a considerable number of ion channels are used. When the trigger signal is received, the gated ion channels in the membrane’s wall open quickly and only briefly. The rushed-in ions create a large offset potential in the layer proximate to the membrane’s surface (from an electrical point of view, a condenser) that drives an intense current through the always-open ion channel array AIS (from an electrical point of view, a resistor) at the beginning of the axon. The slow current on the membrane’s surface mimics a condenser: in the picture of classical electricity, the charges stay inside the distributed condenser while they travel on its surface. In classical electronics, the condenser is sizeless.
7 Thermodynamics of Action Potential
7.1 Deriving thermodynamic description
As the middle inset in Fig. 13 shows, the sum of the rush-in and AIS gradients (plus more gradients if any) controls the generation of the AP. During excitation, $Q_{X}$ energy is fed into the neuron. The rush-in of the $Na^{+}$ ions injects more $Q_{R}$ energy, which can be used to generate an AP. It is a kind of potential energy, which is converted in the first phase of $AP$ (while the voltage gradient is positive) to the kinetic energy $Q_{H}$ of AP. When the membrane potential reaches its peak, the gradient reverses, decreasing the potential. In the second phase of AP (while the voltage gradient is negative) the neuron harvests $Q_{L}$ energy. At the beginning of the transient state, the size of the neuron (due to the increased electrostatic repulsion) slightly increases [80] and stores the energy mainly in the form of elastic energy. The resultant of the electrical, elastic, and thermodynamic forces moves the ions toward the AIS. Although part of the energy is dissipated while the ions exit through the AIS with their Stokes-Einstein speed, the rest of the energy (the elastic part) is recovered. The concentration and potential values change as discussed in connection with Fig. 10. The membrane voltage controls synapses as shown in the top inset in Fig. 13.
@@ 7.1 Deriving thermodynamic description
Figure 16: The time course of the electrical gradient [53] perfectly matches that of the mechanical and thermal changes [35]; however, classic theory cannot explain the connection between them [31]. The experimental $\frac{dV}{dt}$ uses an opposite sign convention.
@@
The ‘phase loop’ (how the voltage gradient and membrane voltage correlate; a parametric curve that uses time $t$ as parameter) shown on the left side of Fig. 15 represents a strong test of the theoretical line shape: the values of the potential and its gradient are simultaneously compared to the experimentally measured ones. Comparing the experimental and theoretical shapes supports the theory that describes the mutual dependence of voltage and its gradient. Furthermore, it shows that omitting the gradient from the classic theory misled the experiment designers, and the measurement design was wrong: the time interval between recording data points was too long. Where the gradient is high and changes quickly (see Fig. 13, middle inset), the blue experimental diagram line is roughly broken. The changes in the values are too significant, causing the integration to distort the value and resulting in an inaccurate broken line. In the rush-in region, the theoretical line is even outside the figure’s range. Where the gradient change is slow, the measurement points are sufficiently dense, and the theory perfectly describes the experimental data. Neither of the mentioned competing (or other classical) theories can interpret the phenomenon and produce any similar dependence.
@@
An exciting aspect is that the phase plot in Fig. 15, left side, shows a remarkable resemblance to a specialized thermodynamic Carnot cycle in Fig. 15, right side; and inspires us to interpret the parameters of the "neuronal Carnot machine" (the bent arcs serve only for illustration; they do not follow the path on the left side). As discussed in [35], a neuron emits and absorbs heat (energy) in the first and second phases, respectively, of an AP; see Fig. 16. The experience, alone, defies the classical HH mechanism: the leakage current can only emit heat, but not absorb it. Laws describing currents of electrons must not be used directly to currents of ions. Neglecting the repelling force between the charge carriers of the ion current, which can be omitted in the case of electron current, leads to entirely wrong conclusions, from the wrong oscillator model to losing the connection between the electrical and mechanical changes, to losing the connection between inanimate and living nature. Our theoretical model, unlike the others, can explain that observation quantitatively.
@@
We rescale the horizontal axis by the resistance $R$ of the AIS (actually, the horizontal axis is transformed to $I_{AIS}$); thus, the plot becomes a neuronal energy density diagram. Actually, it corresponds to a usual PVP-V thermodynamic diagram, and we expect that the temperature changes in the same analogy (we keep the convention of using the letter $V$ to denote both potential and volume). By integrating the parametric $t$ values of any two points on the diagram ‘circle’ line, we receive the work performed by the neuron.
@@
We consider that the environment in the period ‘Excite’ invests energy $Q_{X}$ through the neuron’s synapses (increases the membrane’s voltage to the threshold) between [0,1] to open the valves. Opening the valves enables the neuron receiving in period ‘Rushin’ energy $Q_{R}$ by letting in $Na^{+}$ ions, accelerated by the field across the membrane. Suddenly, the ion concentration and the membrane potential reach their maximum values (given that the current is slow, the sudden change generates a smeared and delayed current on the AIS, see Fig. 13, middle inset). Given that the process is instant, the potential $V$ does not change (no current flows), so the integral [1,2] evaluates to zero (an ‘iso-voltage’ transition). Now the neuron has a high potential to prepare its (already encoded) message (the AP). In terms of electricity, in the first phase (from the threshold potential to the maximum value of the AP, where $\frac{dV}{dt}$ is positive), the neuronal power cycle produces a positive heat $Q_{H}$, given that $\frac{dV}{dt}$ is always positive. In the second phase (from the top value of AP to where the driving force $\frac{dV}{dt}$ becomes zero again), the neuronal power cycle produces a negative heat $Q_{L}$, given that $\frac{dV}{dt}$ is always negative; in line with the measurement results shown in Fig. 16 [35]. The integral [2,3] provides the “heating” $Q_{H}$, the energy invested. Similarly, the integral [3,6] provides the “cooling” $Q_{L}$, representing the energy recovered (assuming no artificial current).
In terms of thermodynamics, the ions stop in the immediate vicinity of the membrane, i.e., initially, the pressure created by their mutual repulsion is present at the inner surface of the sphere (it is a mechanical shock wave). In the positive phase of the damped oscillation, it compresses the fluid in the sphere. In the second phase, in the negative phase of oscillation, it decompresses the fluid (sucks back the fluid). Given that the density of ions in the volume is constant (the ions’ speed is too low to follow those sudden changes), the electrical and thermodynamic quantities faithfully follow each other. Measuring electrical potential and mechanical pressure yields disciplinary measurements of the same physical effect.
In the “Relax” phase, $\frac{dV}{dt}$ is (almost) zero, so the corresponding integral evaluates to zero (‘iso-gradient’ transition). In this phase, the neuron must restore its resting state: by using ATP, it creates new ions. The difference in the energies of $Q_{H}$ and $Q_{L}$ is the net energy used to forward an AP; while $Q_{X}$ is the energy used to perform the computation. The energy $Q_{X}$ is an approximate value: as shown in Fig. 13, at the time of re-opening the synaptic inputs, the membrane can be above or below the resting potential, which actually means a neuron-level memory. In this sense, adjacent neural spikes can borrow energy from each other (which explains why the spikes in a burst behave differently). The energy $Q_{R}$ is produced by the ATP hydrolysis in parallel with the ’primary processes’ described above, and the produced ions are acquired on the condenser plates, thus restoring the resting potential and providing potential energy for the forthcoming spikes.
@@
From Fig. 15, after recalibration (with $R_{AIS}=10^{7}$ Ohm), one can estimate (by approximating roughly the area under the arcs by triangles, in units of $10^{-7}[J]$), $Q_{H}=5.2,Q_{L}=1.8,W=3.4,Q_{X}=0.2$. That is, according to the "phase plot" [53], the neuron operates with an efficiency around 65%, and approximately 3% of the energy is derived from the upstream neurons. The values (without using data from dedicated measurements) agree excellently with the measured [43] values $W=2.5$ and 74%; furthermore, the $Q_{X}$:total ratio $1:35$, derived by measuring the direct energy consumption [45]. An interesting proof that the two disciplines see the same physical effect, that electrical parameters can measure the thermodynamic efficiency. The invested energy can be safely estimated from the work the rush-in charge performs given that the energy is almost entirely stored as elastic potential energy due to the enormous elasticity modulus